# Radial Transport in High-Redshift Disk Galaxies Dominated by Inflowing Streams

### v1

@@ 1 Introduction
Extended, self-gravitating, star-forming, highly turbulent, rotating disks dominate galaxy formation at cosmic noon, $z\sim 3-1$ (e.g., Erb et al., 2004; Förster Schreiber et al., 2006; Genzel et al., 2006; Elmegreen et al., 2007; Kassin et al., 2007; Genzel et al., 2008; Stark et al., 2008; Förster Schreiber et al., 2009; Kassin et al., 2012; Swinbank et al., 2012; Wisnioski et al., 2015; Stott et al., 2016; Simons et al., 2017; Swinbank et al., 2017; Übler et al., 2019; Wisnioski et al., 2019), representing the peak of star formation activity in the Universe (Madau & Dickinson, 2014).
@@
Another mechanism is turbulent viscosity. According to Krumholz & Burkert (2010), turbulence can be modeled as an effective viscosity, which induces torques in a marginally unstable disk. The gravitational energy gained from the resulting inward mass transport balances the energy lost due to dissipation of turbulence such that an energy equilibrium is maintained. Krumholz et al. (2018) extend the Krumholz & Burkert (2010) model, considering turbulence driven by both disk instability and supernova feedback, where energy lost due to dissipation of turbulence is balanced by both energy gained from supernova feedback and VDI-driven radial mass transport in a marginally unstable disk. They find that while supernova-feedback-driven turbulence is sufficient to explain the observed velocity dispersions in low mass galaxies at low redshifts, disk-instability-driven turbulence is required at high redshifts (see also Forbes et al., 2012, 2014). The Krumholz et al. (2018) model is extended by Ginzburg et al. (2022), who include an additional energy input from external accretion and show that while supernova feedback is the main driver of turbulence in low mass disks (halo masses $<10^{11.5}M_{\odot}$ at $z\sim 2$), the primary driver of turbulence in high mass disks is either external accretion or VDI-based radial mass transport, depending on the efficiency by which stream kinetic energy is converted to turbulent energy, which is treated as a free parameter (see also Elmegreen & Burkert, 2010; Klessen & Hennebelle, 2010; Genel et al., 2012a).
@@
We recall that disks at cosmic noon are fed by cold gas streams from the cosmic web (Birnboim & Dekel, 2003; Kereš et al., 2005; Dekel & Birnboim, 2006; Ocvirk et al., 2008; Dekel et al., 2009a; Danovich et al., 2012, 2015; Mandelker et al., 2016; Padnos et al., 2018; Mandelker et al., 2018; Aung et al., 2019; Mandelker et al., 2019, 2020). The streams are likely to play an additional role in driving the radial transport of cold gas in disks, either directly by their inward velocities and/or indirectly by driving disk turbulence. This is expected to introduce deviations from cylindrical symmetry, which are not captured by the VDI models. Investigating the radial transport of cold gas in $z\sim 0$ Milky Way mass disks using zoom-in hydro-cosmological simulations of the FIRE suite (Hopkins et al., 2014), Trapp et al. (2022) find that most of the freshly incoming gas piles up near the disk edge and does not contribute much to the radial mass flux within the disk. However, the situation in cosmic-noon disks may be different, as they are fed by more intense streams (Dekel & Birnboim, 2006; Dekel et al., 2009a). Using the VELA simulations, we measure the direct contribution of the streams to the radial transport of cold gas in cosmic-noon disks. Distinguishing freshly incoming streams from the off-stream disk gas is non-trivial in an Eulerian simulation, and we resort to crude stream identification techniques (see §5 for details).
@@ 2.1 VELA Simulations
The VELA simulation suite consists of 34 hydro-cosmological zoom-in simulations of galaxies, typically evolved to $z\sim 1$, with halo masses in the range of $\sim\!10^{11}-10^{12}\>{\rm M_{\odot}}$ at $z\sim 2$ (Ceverino et al., 2014; Zolotov et al., 2015; Ceverino et al., 2016a, b; Mandelker et al., 2017). The simulations are run with the Adaptive Refinement Tree (ART) code (Kravtsov et al., 1997; Kravtsov, 2003; Ceverino & Klypin, 2009), which accurately follows the evolution of a gravitating $N$-body system and Eulerian gas dynamics using an AMR scheme. The maximum AMR resolution of the simulations is $17.5-35\>{\rm pc}$ in physical units at all times. In addition to gravity and hydrodynamics, the simulations incorporate the physics of gas and metal cooling, UV-background photoionization, stochastic star formation, gas recycling, stellar winds and metal enrichment, thermal feedback from supernovae (Ceverino et al., 2010, 2012), and feedback from radiation pressure (Ceverino et al., 2014). For further details on the simulation method and the VELA simulation suite in general, see Ceverino et al. (2010, 2014); Zolotov et al. (2015); Mandelker et al. (2017).
@@ 2.2 Physical Quantities
Except when otherwise stated, here onwards, we focus on the cold gas within the disk, i.e., gas with temperatures less than $1.5\times 10^{4}\>{\rm K}$ and residing within a cylinder of radius $R_{\rm d}$ and height $H_{\rm d}$. We have verified that increasing the temperature threshold of cold gas to as high as $5\times 10^{4}\>{\rm K}$ has no significant impact on our results. The ratio of the cold gas mass to the baryonic (gas + stars) mass within the disk is defined as the cold gas fraction, $f$. In §5, the cold gas within the disk is further classified into recently accreted streams versus the off-stream material, labeled as non-streams.
@@ 2.3 Disk Galaxy Sample
Figure 1: Disk Selection: The ratio of the disk-averaged rotational velocity to the disk-averaged radial velocity dispersion, $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$, is plotted against the ratio of the disk radius to the disk height, $R_{\rm d}/H_{\rm d}$, for all 1098 snapshots belonging to 34 VELA galaxies with the blue, round points. From these snapshots, rotation-supported disks are selected by requiring that $R_{\rm d}/H_{\rm d}>3$ and $V_{\rm rot}/\sigma_{r}>1$ at all $r$, the distance from the galactic center in cylindrical coordinates. Here, $V_{\rm rot}$ and $\sigma_{r}$ are the average rotational velocity and the radial velocity dispersion of cold gas as a function of $r$, respectively. The red stars highlight the (selected) disk galaxy sample and comprise 425 snapshots, about $\sim 39\%$ of all VELA snapshots. The gray, dashed vertical and horizontal lines denote $R_{{\rm d}}/H_{{\rm d}}=3$ and $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle=1$, respectively. Note the blue points with $R_{\rm d}/H_{\rm d}>3$ and $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle>1$ that are not in the disk galaxy sample, as they do not satisfy the condition $V_{\rm rot}/\sigma_{r}>1$ at all $r$. However, including these snapshots in the disk sample has no significant effect on our results.
@@
Figure 2: Radial Transport in a VELA Disk: For the simulated galaxy VELA 7 at a redshift of $z=1.5$, the top left- and right-hand panels, respectively, show the maps of face-on projected surface density, $\Sigma$, and radial velocity, $v_{r,\rm 2D}$, for the cold gas in a region of dimensions $3\ R_{\rm d}\times 3\ R_{\rm d}$, centered on the galactic center. The disk is large ($R_{\rm d}=18.43\ \rm kpc$) and thin ($R_{\rm d}/H_{\rm d}\sim 9$). It has a dominant rotational support with $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle\sim 3$ and a gas fraction of $0.18$. Regions devoid of cold gas are shown with white and black patches in the surface density and radial velocity maps, respectively. The red, dashed circles indicate the radial extent of the disk. Note that both the radial velocity and surface density maps are highly non-axisymmetric, largely representing incoming streams of cold gas. The bottom left- and right-hand panels, respectively, show the average radial velocity, $V_{r}$ (in units of $V_{\rm rot}$), and radial mass flux, $F_{r}$ (in units of $\mathscr{M}/t_{\rm dyn}$), as functions of $r$ (in units of $R_{\rm d}$) out to the disk radius. Here, $V_{\rm rot}$ and $\mathscr{M}$ are the average rotational velocity and mass of all cold gas at $r$, and $t_{\rm dyn}=r/V_{\rm rot}$ is the local dynamical time. Except within $\sim\!0.2\ R_{\rm d}$, $V_{r}$ and $F_{r}$ are both negative and generally increase in magnitude with an increase in $r$, though there are some oscillations. The disk-averaged radial velocity is also negative and about $8\%$ of the disk-averaged rotational velocity in magnitude. Therefore, the radial transport of cold gas in the disk is inflow-dominated.
@@
Figure 3: Trends with Distance from the Galactic Center: The left- and right-hand panels, respectively, show the medians (solid lines) and $16^{\rm th}-84^{\rm th}$ percentiles (envelopes) in the average radial velocity, $V_{r}$ (in units of $V_{\rm rot}$), and radial mass flux, $F_{r}$ (in units of $\mathscr{M}/t_{\rm dyn}$), as functions of $r$ (in units of $R_{\rm d}$) obtained from our disk galaxy sample. The median $V_{r}$ and $F_{r}$ are always negative and increase in magnitude with an increase in $r$, except within $\sim\!0.2\ R_{\rm d}$, where the trends are reversed. While both $F_{r}$ and $V_{r}$ are positive at the $84^{\rm th}$ percentile for all $r$, their magnitudes at the $16^{\rm th}$ percentile are much larger. Therefore, in a statistical sense, the radial transport of cold gas in our disk galaxy sample is inflow-dominated at all radii.
@@
From the 1098 snapshots, rotation-supported disks are selected by requiring that $R_{\rm d}/H_{\rm d}>3$ and $V_{\rm rot}/\sigma_{r}>1$ at all $r$. The former ensures a disk-like geometry, while the latter affirms a rotational support that is larger than the dispersion support at all $r$. The (selected) disk galaxy sample comprises 425 snapshots, about $\sim 39\%$ of all VELA snapshots. In Fig. 1, $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$, which is a measure of the disk-averaged rotational support in the galaxy compared to the disk-averaged dispersion support, is plotted against $R_{\rm d}/H_{\rm d}$ for all VELA snapshots with the blue, round points. The red stars highlight the disk galaxy sample. The gray, dashed vertical and horizontal lines denote $R_{\rm d}/H_{\rm d}=3$ and $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle=1$, respectively. Note the blue points with $R_{\rm d}/H_{\rm d}>3$ and $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle>1$ that are not in the disk galaxy sample, as they do not satisfy the condition $V_{\rm rot}/\sigma_{r}>1$ at all $r$. However, we have verified that including these snapshots in the disk sample makes no significant difference to our results.
@@ 3.1 An Illustrative Example
Figure 4: Trends with Cold Gas Fraction and Redshift: The left- and right-hand panels, respectively, plot the ratio of the disk-averaged radial velocity to the disk-averaged rotational velocity, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$, against the cold gas fraction, $f$, and redshift, $z$, with the blue, round points. The red squares and the vertical error bars denote the medians and $16^{\rm th}-84^{\rm th}$ percentiles, respectively, in $f$ and $z$ bins of varying widths (to include a sensible number of snapshots), as indicated by the horizontal error bars, with the locations of the squares along the $f$- and $z$-axes being at the bin-averages. For most of the VELA disks ($82\%$), $\langle V_{r}\rangle$ is negative, highlighting the inflow-dominated nature of cold gas radial transport in our disk galaxy sample (note that $\langle V_{\rm rot}\rangle$ is always positive). Also, on average, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ decreases (increases in magnitude) with an increase in both $f$ and $z$, indicating that galaxies with larger cold gas fractions or at higher redshifts have higher levels of disk-averaged radial inflow in general. Both trends are rather weak, though the one with cold gas fraction is more pronounced than that with redshift, the Spearman’s rank correlation coefficients in the two cases being $-0.26$ and $-0.13$, respectively.
@@
In the bottom left- and right-hand panels, we, respectively, show the cold gas $V_{r}$ (in units of $V_{\rm rot}$) and $F_{r}$ (in units of $\mathscr{M}/t_{\rm dyn}$) as functions of $r$ (in units of $R_{\rm d}$) out to the disk radius. Except within $\sim\!0.2\ R_{\rm d}$, $V_{r}$ and $F_{r}$ are both negative and generally increase in magnitude with an increase in $r$, though there are some oscillations. $V_{r}$ is most negative near $r=0.8\ R_{\rm d}$, attaining a value of $-0.2\ V_{\rm rot}$. $F_{r}$ is also most negative at the same radius. The disk-averaged radial velocity, $\langle V_{r}\rangle$, is also negative and about $8\%$ of the disk-averaged rotational velocity, $\langle V_{\rm rot}\rangle$, in magnitude. Therefore, the radial transport of cold gas in the disk is inflow-dominated.

### v2

# Radial Transport in High Redshift Disk Galaxies Dominated by Inflowing Streams

@@ 1 Introduction
Extended, self-gravitating, star-forming, highly turbulent, rotating disks dominate galaxy formation at cosmic noon, $z\sim 3-1$ (e.g., Erb et al., 2004; Förster Schreiber et al., 2006; Genzel et al., 2006; Elmegreen et al., 2007; Kassin et al., 2007; Genzel et al., 2008; Stark et al., 2008; Förster Schreiber et al., 2009; Kassin et al., 2012; Swinbank et al., 2012; Wisnioski et al., 2015; Stott et al., 2016; Simons et al., 2017; Swinbank et al., 2017; Übler et al., 2019; Wisnioski et al., 2019), representing the peak of the star formation rate density in the Universe (Madau & Dickinson, 2014).
@@
Another mechanism is turbulent viscosity. According to Krumholz & Burkert (2010), turbulence can be modeled as an effective viscosity, which induces torques in a marginally unstable disk. The gravitational energy gained from the resulting inward mass transport balances the energy lost due to dissipation of turbulence such that an energy equilibrium is maintained (see also Goldbaum et al., 2015). Krumholz et al. (2018) extend the Krumholz & Burkert (2010) model, considering turbulence driven by both disk instability and supernova feedback, where energy lost due to dissipation of turbulence is balanced by both energy gained from supernova feedback and VDI-driven radial mass transport in a marginally unstable disk. They find that while supernova-feedback-driven turbulence is sufficient to explain the observed velocity dispersions in low mass galaxies at low redshifts, disk-instability-driven turbulence is required at high redshifts (see also Forbes et al., 2012, 2014; Goldbaum et al., 2016). The Krumholz et al. (2018) model is extended by Ginzburg et al. (2022), who include an additional energy input from external accretion and show that while supernova feedback is the main driver of turbulence in low mass disks (halo masses $10^{11.5}M_{\odot}$ at $z\sim 2$), the primary driver of turbulence in high mass disks is either external accretion or VDI-based radial mass transport, depending on the efficiency by which stream kinetic energy is converted to turbulent energy, which is treated as a free parameter (see also Elmegreen & Burkert, 2010; Klessen & Hennebelle, 2010; Genel et al., 2012a; Forbes et al., 2023).
@@
We recall that in massive halos at cosmic noon, cold gas from the cosmic web is funneled into hot halos through a few dominant filaments or streams, delivering it deep within the halo to the host galaxy (Birnboim & Dekel, 2003; Kereš et al., 2005; Dekel & Birnboim, 2006; Ocvirk et al., 2008; Dekel et al., 2009a; Pichon et al., 2011; van de Voort et al., 2011; Danovich et al., 2012, 2015; Nelson et al., 2013, 2015; Padnos et al., 2018; Mandelker et al., 2018; Aung et al., 2019; Mandelker et al., 2019, 2020). These streams are likely to play an additional role in driving the radial transport of cold gas in disks, either directly by their inward velocities and/or indirectly by driving disk turbulence. This is expected to introduce deviations from cylindrical symmetry, which are not captured by the VDI models.
[…]
@@ 2.1 VELA Simulations
The VELA simulation suite consists of 34 hydro-cosmological zoom-in simulations of galaxies, typically evolved to $z\sim 1$, with halo masses in the range of $\sim\!10^{11}-10^{12}\>{\rm M_{\odot}}$ at $z\sim 2$ (Ceverino et al., 2014; Zolotov et al., 2015; Ceverino et al., 2016a, b; Mandelker et al., 2017). The simulations are run with the Adaptive Refinement Tree (ART) code (Kravtsov et al., 1997; Kravtsov, 2003; Ceverino & Klypin, 2009), which accurately follows the evolution of a gravitating $N$-body system and Eulerian gas dynamics using an AMR scheme. The maximum AMR resolution of the simulations is $17.5-35\>{\rm pc}$ in physical units at all times. In addition to gravity and hydrodynamics, the simulations incorporate the physics of gas and metal cooling, UV-background photoionization, stochastic star formation, gas recycling, stellar winds and metal enrichment, thermal feedback from supernovae (Ceverino et al., 2010, 2012), and feedback from radiation pressure (Ceverino et al., 2014). The dark matter particles have a mass of $8.3\times 10^{4}M_{\odot}$ and stellar particles are formed with a minimum mass of $10^{3}M_{\odot}$. For further details on the simulation method and the VELA simulation suite in general, see Ceverino et al. (2010, 2014); Zolotov et al. (2015); Mandelker et al. (2017).
@@ 2.2 Physical Quantities
[…]
@@
Figure 1: Disk Selection: The ratio of the disk-averaged rotational velocity to the disk-averaged radial velocity dispersion, $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$, is plotted against the ratio of the disk radius to the disk height, $R_{\rm d}/H_{\rm d}$, for all 1098 snapshots belonging to 34 VELA galaxies with the blue, round points. Of these, 425 snapshots satisfy the conditions $R_{\rm d}/H_{\rm d}>3$ and $V_{\rm rot}/\sigma_{r}>1$ at all $r$, the distance from the galactic center in cylindrical coordinates, and are selected as rotation-supported disks, highlighted by the red stars. Here, $V_{\rm rot}$ and $\sigma_{r}$ are the average rotational velocity and the radial velocity dispersion of cold gas as a function of $r$, respectively. The gray, dashed vertical and horizontal lines indicate $R_{{\rm d}}/H_{{\rm d}}=3$ and $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle=1$, respectively. Some of the blue points with $R_{\rm d}/H_{\rm d}>3$ and $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle>1$ do not satisfy the condition $V_{\rm rot}/\sigma_{r}>1$ at all $r$. As such, they are not in the disk galaxy sample.
@@ 2.3 Disk Galaxy Sample
Figure 2: Radial Transport in a VELA Disk: For the simulated galaxy VELA 7 at a redshift of $z=1.5$, the top left- and right-hand panels show face-on maps of the cold gas surface density ($\Sigma$) and the two-dimensional radial velocity ($v_{r,\rm 2D}$), respectively, within a region spanning $3\ R_{\rm d}\times 3\ R_{\rm d}$, centered on the galactic center. The disk is large ($R_{\rm d}=18.43\ \rm kpc$) and thin ($R_{\rm d}/H_{\rm d}\sim 9$). It is mainly supported by rotation, with $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle\sim 3$, and has a cold gas fraction of $f=0.18$. Regions with no cold gas are shown as white in the surface density map and black in the radial velocity map. Dashed red circles mark the radial extent of the disk. Both maps exhibit strong non-axisymmetry, largely due to incoming cold gas streams. The middle left- and right-hand panels display the average radial velocity ($V_{r}$, in units of $V_{\rm rot}$) and the radial mass flux ($F_{r}$, in units of $\mathscr{M}/t_{\rm dyn}$), respectively, as functions of $r$, normalized by $R_{\rm d}$, out to the disk radius. Here, $V_{\rm rot}$ and $\mathscr{M}$ are the average rotational velocity and mass of all cold gas at $r$, and $t_{\rm dyn}=r/V_{\rm rot}$ is the local dynamical time. The non-normalized $V_{r}$ and $F_{r}$ are shown in the bottom left- and right-hand panels, respectively. Beyond roughly $0.2\ R_{\rm d}$, both $V_{r}$ and $V_{\rm rot}$ are negative and tend to grow in magnitude with increasing $r$, albeit with some fluctuations, indicating a predominantly inflow-driven cold gas radial transport within the disk. The disk-averaged radial velocity is also negative and has a magnitude about $8\%$ of the disk-averaged rotational velocity.
@@
Figure 3: Trends with Distance from the Galactic Center: The left- and right-hand panels show, respectively, the medians (solid curves) and the $16^{\rm th}$$84^{\rm th}$ percentile ranges (shaded regions) for the average radial velocity ( $V_{r}$, in units of $V_{\rm rot}$), and the radial mass flux ($F_{r}$, in units of $\mathscr{M}/t_{\rm dyn}$), as functions of $r$ (in units of $R_{\rm d}$), derived from our sample of disk galaxies. The median values of both $V_{r}$ and $F_{r}$ are consistently negative and become more negative with increasing $r$, except within approximately $\sim\!0.2\ R_{\rm d}$, where the trend reverses. Although the $84^{\rm th}$ percentile values of both $F_{r}$ and $V_{r}$ are positive at all radii, the magnitudes at the $16^{\rm th}$ percentile are substantially larger. Therefore, statistically speaking, the radial transport of cold gas across our sample is dominated by inflows at all radii.
@@
From the 1098 snapshots, rotation-supported disks are selected by requiring that $R_{\rm d}/H_{\rm d}>3$ and $V_{\rm rot}/\sigma_{r}>1$ at all $r$. The former is a reasonable cut to ensure a disk-like geometry, while the latter affirms a rotational support that is larger than the dispersion support at all $r$. This selection is both for comparison to disk-instability-based transport models, which are valid for rotation-supported disks, and also to the Genzel et al. (2023) sample of galaxies. The minimum $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ of our (selected) disk galaxy sample is around $1.5$, while that of the Genzel et al. (2023) galaxies is $2$. Our sample comprises 425 snapshots, about $\sim 39\%$ of all VELA snapshots. In Fig. 1, $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$, which is a measure of the disk-averaged rotational support in the galaxy compared to the disk-averaged dispersion support, is plotted against $R_{\rm d}/H_{\rm d}$ for all VELA snapshots with the blue, round points. The red stars highlight the disk galaxy sample. The gray, dashed vertical and horizontal lines denote $R_{\rm d}/H_{\rm d}=3$ and $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle=1$, respectively. Note the blue points with $R_{\rm d}/H_{\rm d}>3$ and $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle>1$ that are not in the disk galaxy sample, as they do not satisfy the condition $V_{\rm rot}/\sigma_{r}>1$ at all $r$. However, we have verified that including these snapshots in the disk sample makes no significant difference to our results.
@@ 3.1 An Illustrative Example
Figure 4: Trends with Cold Gas Fraction and Redshift: The left- and right-hand panels show the ratio of the disk-averaged radial velocity to the disk-averaged rotational velocity, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$, plotted as a function of the cold gas fraction ($f$) and redshift ($z$), respectively, using blue circles. Red squares represent the median values within bins of $f$ and $z$ (chosen to ensure adequate snapshot counts), while the vertical error bars correspond to the $16^{\rm th}$$84^{\rm th}$ percentile ranges. The horizontal error bars indicate the bin widths, and the squares are placed at the mean values of $f$ and $z$ for each bin. For the majority of VELA disks ($82\%$), $\langle V_{r}\rangle$ is negative (note that $\langle V_{\rm rot}\rangle$ is strictly positive), underscoring the dominance of cold gas radial inflows over outflows in our disk galaxy sample. On average, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ becomes more negative with increasing $f$ and $z$, suggesting that galaxies with higher cold gas content or located at higher redshifts typically exhibit stronger radial inflows. Although both correlations are relatively weak, the trend with cold gas fraction is stronger than that with redshift, as reflected in the Spearman rank correlation coefficients of 0.26-0.26 and 0.13-0.13, respectively.
@@
In the middle left- and right-hand panels, we, respectively, show the cold gas $V_{r}$ (in units of $V_{\rm rot}$) and $F_{r}$ (in units of $\mathscr{M}/t_{\rm dyn}$) as functions of $r$ (in units of $R_{\rm d}$) out to the disk radius. The non-normalized $V_{r}$ and $F_{r}$ are depicted in the bottom left- and right-hand panels, respectively. Except within $\sim\!0.2\ R_{\rm d}$, $V_{r}$ and $F_{r}$ are both negative and generally increase in magnitude with an increase in $r$, though there are some oscillations. $V_{r}$ is most negative near $r=0.8\ R_{\rm d}$, attaining a value of $0.2\ V_{\rm rot}$. $F_{r}$ is also most negative at the same radius. The disk-averaged radial velocity, $\langle V_{r}\rangle$, is also negative and about $8\%$ of the disk-averaged rotational velocity, $\langle V_{\rm rot}\rangle$, in magnitude. Therefore, the radial transport of cold gas in the disk is inflow-dominated.